# Can Machines Resonate with Humans? Evaluating the Emotional and Empathic Comprehension of LMs

### v1

@@ Abstract
Empathy plays a pivotal role in fostering prosocial behavior, often triggered by the sharing of personal experiences through narratives. However, modeling empathy using NLP approaches remains challenging due to its deep interconnection with human interaction dynamics. Previous approaches, which involve fine-tuning language models (LMs) on human-annotated empathic datasets, have had limited success. In our pursuit of improving empathy understanding in LMs, we propose several strategies, including contrastive learning with masked LMs and supervised fine-tuning with Large Language Models (LLMs). While these methods show improvements over previous methods, the overall results remain unsatisfactory. To better understand this trend, we performed an analysis which reveals a low agreement among annotators. This lack of consensus hinders training and highlights the subjective nature of the task. We also explore the cultural impact on annotations. To study this, we meticulously collected story pairs in Urdu language and find that subjectivity in interpreting empathy among annotators appears to be independent of cultural background. The insights from our systematic exploration of LMs’ understanding of empathy suggest that there is considerable room for exploration in both task formulation and modeling.
@@ 1 Introduction
With Large Language Models (LLMs) demonstrating impressive capabilities in generating naturally sounding answers over a broad range of human inquiries, more individuals turn to seek solutions and emotional comfort by interacting with LLM-supported chatbots OpenAI (2023); Chang et al. (2024). They express thoughts, feelings and share their experiences, expecting deep understanding and sympathetic responses from a chatbot that can resonate with them, as shown in Figure 1 (Berridge et al., 2023). This requests LLMs to first fully understand the event, the emotions, and the empathy in the narratives, and then to respond appropriately Lin et al. (2023).
@@
Shen et al. (2023) proposed the task of measuring empathic similarity. It assesses the similarity between two narratives describing personal experiences from four aspects including event, emotion, moral, and empathy by a numerical score ranging from 1 to 4. However, both fine-tuned LMs (BERT, RoBERTa) and few-shot prompting LLMs achieved low correlation with human annotations. Our error analysis shows that models can recognize dissimilar story pairs with scores in range of 1-2, but struggle to distinguish fine-grained differences between similar story pairs in range of 2.5-4 (see Section 2). This suggests that the nuanced patterns and relationships in similar pairs are not captured by current methods Wang et al. (2022).
We hypothesize that the LMs used in these methods were primarily trained for semantic understanding tasks, rather than for emotion and empathy. They may capture the similarity between events, but they struggle to understand the complex social and emotional signals in human narratives Reimers and Gurevych (2019), this is further discussed in Appendix A. Moreover, naïve LLM prompts may not fully empower LLMs’ reasoning ability to make correct judgements. To this end, we explore multiple strategies to improve the accuracy of empathic similarity, including contrastive learning of LMs, LLM reasoning and fine-tuning with and without Chain-of-Thought (CoT) (Wei et al., 2022). We achieve around 5–10% improvement in the Pearson and Spearman correlation. However, accuracy appears to be capped and cannot be improved beyond a certain value.
This motivated us to speculate that gold labels might be problematic due to the subjective nature of judging emotion and empathy similarity. To investigate this, we randomly sampled ten story pairs from EmpathicStories and asked eight annotators with different background to annotate the similarity in terms of event, emotion, moral and empathy. Overall low human agreement for all aspects was observed, especially moral, followed by empathy, emotion; the highest agreement was for the event. Interestingly, annotators from the same country/culture, particularly who are friends, had much higher agreement than others. We further collected a new Urdu dataset to explore the cultural impact on empathy similarity labeling, which revealed similar findings.
@@
  • We collect a new Urdu dataset to investigate the impact of language and culture on empathic labeling. Our analysis concludes that subjectivity in the labeling is independent of the cultural background of annotators.
@@ 2.1 Task and Dataset
In this work, we focus on the task of measuring the empathic similarity between two narratives by a numerical score in the range of 1 to 4, with 1 representing totally dissimilar and 4 indicating extremely similar. We used the EmpathicStories dataset, which was created by Shen et al. (2023). It consists of 1,500 unique stories and 2,000 story pairs, split into 1,500 pairs for training, 100 for development, and 400 for testing. Table 1 shows the distribution of train, development, and test sets across different similarity ranges. Each story has two versions, full and summary, in which the annotators assign labels based on the summary from four perspectives including event, emotion, moral, and overall empathy. Figure 2 shows the correlation between the four similarity scores. We can see that the moral similarity has the highest correlation with empathy, followed by event and emotion.
Evaluation Measurements
Pearson correlation ($r$) and Spearman rank correlation ($\rho$) are used to measure the performance of systems for predicting emotional similarity. This measure the linear correlation between the model outputs and the human annotations, the degree of monotonicity under ranking, respectively. Mean square error (MSE) assesses the models’ ability to get close to the gold standard. In a discrete setting, following Shen et al. (2023), both the predicted scores and the gold labels are binned into two classes thresholding by 2.5 — label = 1 if score > 2.5; otherwise, it is 0. Then accuracy, precision, recall and macro-F1 are used.
@@ 2.2 Error Analysis
Baseline: we reproduced the results of Shen et al. (2023) by fine-tuning SBERT Reimers and Gurevych (2019) and BART-base Lewis et al. (2019) as Table 9, given that prompting davinci-text-003 and GPT-3.5-turbo had lower accuracy. We obtained Spearman correlation $\rho$ that was similar to that in the original paper, with lower F1.
@@
We found that the model excels at identifying dissimilar story pairs, but struggles with similar pairs with nuanced differences. Specifically, the model exhibits lower error rates in the score range of 1–2, and higher error rates in the range of 2.5–4. Particularly for 4, both models have a 100% error rate on the test set. We could attribute this to exposure bias during training. However, the number of training examples scoring between 2.5 and 4 is actually much larger than the number of dissimilar examples. This suggests that it is difficult for models to discern the subtle differences between similar story pairs, underlining a critical area for model improvement.
Hypothesis
We hypothesize that SBERT and BART are primarily pre-trained and fine-tuned to learn semantic features, in which emotional and empathetic features are under-represented. Despite the importance of understanding the semantics of the stories, this does not necessarily account for capturing the deep, empathetic connections between narratives, especially when these connections are subtler and more implicit than straightforward lexical or thematic commonalities. The failure to identify the connections between similar narratives in terms of empathy leads to poor performance on similar story pairs ranging from 2.5 to 4, tending to be recognized as dissimilar with a score between 1-2. To alleviate these issues and to enhance the model performance on empathetic similarity task, we propose a variety of strategies in Section 3.
@@ 3 Enhancing Empathy Representation
We explored various contrastive losses, including the ContrastiveLoss (Hadsell et al., 2006), AnglELoss (Li and Li, 2024) and CoSENTLoss222www.sbert.net. We also re-implemented the approach using cosine similarity loss as a baseline. We used masked language models (MLMs) including MPNet (Song et al., 2020), RoBERTa (base, large) (Liu et al., 2019) as text encoders leveraging their bidirectional encoding capabilities and computational efficiency.
LLM Reasoning and Fine-tuning
@@
For the CoT explanation, even though the human-annotated explanations were provided in EmpathicStories, they were collected based on the story summary by GPT-3.5-Turbo, which can be biased and inaccurate. Moreover, the distribution of human-written text is distinct from the distribution of LLM-generated reasons, which makes the reasons provided by the humans less useful for lowering the perplexity in LLM fine-tuning  (Liu et al., 2024a). To address this, we prompted a powerful LLM to generate explanations guided by gold labels, leveraging LLM capability of causal reasoning. That is, to explain why two narratives have the similarity of [scores placeholder] from the aspects of event, emotion, moral, and empathy.
Gold Label-Guided Explanation
Based on the pair of full stories with the ground truth similarity scores, we asked the LLM to analyze the story pair from dimensions such as theme, emotional content, characters, narrative structures, and overall empathy. The generated analysis served as the reasoning content in supervised fine-tuning (SFT). We explored both full parameter and PEFT Hu et al. (2021) fine-tuning.
@@ 4.1 Discriminative LMs
We refer to discriminative models as small pre-trained LMs such as RoBERTa, and sentence embedding models like SBERT. For both the zero-shot and the fine-tuning settings, we calculated the cosine similarity based on embeddings of two stories derived from pre-trained or fine-tuned sentence encoders.
@@ 4.1.2 Fine-Tuning with Contrastive Loss
We examined the effectiveness of contrastive losses presented in Section 3 based on three LMs. We fine-tuned them on the four types of gold labels separately and evaluated them using the overall empathy score as ground-truth based on pairs of story summary. [^3] The full story always exceeds the maximum input length of these three LMs, resulting in poor performance. This allows us to discern the influence of each attribute on the empathy similarity modeling.
@@ 4.2.1 Zero-Shot Evaluation
Table 3 shows two major findings: (i) Zero-shot LLMs have clear advantages over cosine similarity based on sentence embeddings. $r$=0.442 for GPT-4o vs. $r$=0.336 for OpenAI-text-embedding on story summary. The larger the model, the higher the correlations with the gold empathy scores. This suggests that LLMs benefit from large-scale pre-training that empowers the distinguishability of nuanced empathic differences in narratives. (ii)!Over all models and over the four types of scores, using the summaries is superior to using the full stories, which aligns with the fact that the gold labels are annotated based on the summary.
@@ 4.2.2 Is SFT Helpful?
To further investigate the potential of LLMs in this task, we performed supervised fine-tuning (SFT) with LLaMA-3-8B using the story summaries and the corresponding empathic similarity scores from the training set. In our experiments, we explored two different prompting strategies as introduced in Section 3. For the Score-Only strategy, we implemented both full-parameter and LoRA (Hu et al., 2021) SFT. For the CoT strategy, we only performed LoRA SFT due to the memory pressure of the longer sequence length. For all strategies, we tuned the model for two epochs.
@@ 4.3 Understanding the Bottleneck
In order to understand what LLMs learned in the fine-tuning and why they struggle with this task, we conducted an analysis of the predictions made by the full-parameter fine-tuned model. We chose the full-parameter fine-tuned model as it has been adapted to the task of empathy similarity estimation, which reduces the interference from its prior knowledge.
@@
Regardless of the similarity class, the predicted probability kept the same as the empirical distribution of training data over four classes, i.e., $P(Y)$ = (0.140, 0.399, 0.404, 0.057), which is calculated by counting the percentage of story pairs falling into each class. That is, after fine-tuning, the model learned nothing about how to map the input $x\rightarrow y$, but merely the distribution of $P(Y)$. During inference, the model could not estimate the corresponding score conditioned on the input story pair, but sampled a similarity class based on the distribution of $P(Y)$ whatever the input pair was, leading to randomness on the development and the test sets (see Figure 4).
Based on these observations and previous findings in semantic textual similarity that models would be confused when training with ambiguous and subjective labels (Wang et al., 2023, 2024), we speculate that the annotated empathic similarity scores are substantially subjective, i.e., that labels of high human disagreements hinder the model to learn distinguishable patterns. This could also explain why LMs and LLMs could not improve the performance by a large margin despite our extensive efforts. The struggle of the models to overcome the bottleneck can be attributed to the complexity and the variability in the human interpretations of these abstract concepts. This highlights the necessity to critically analyze the quality of the dataset and to address the inherent subjectivity in its labels.

### v2

@@ Abstract
Empathy plays a pivotal role in fostering prosocial behavior, often triggered by the sharing of personal experiences through narratives. However, modeling empathy using NLP approaches remains challenging due to its deep interconnection with human interaction dynamics. Previous approaches, which involve fine-tuning language models (LMs) on human-annotated empathic datasets, have had limited success. In our pursuit of improving empathy understanding in LMs, we propose several strategies, including contrastive learning with masked LMs and supervised fine-tuning with large language models. While these methods show improvements over previous methods, the overall results remain unsatisfactory. To better understand this trend, we performed an analysis which reveals a low agreement among annotators. This lack of consensus hinders training and highlights the subjective nature of the task. We also explore the cultural impact on annotations. To study this, we meticulously collected story pairs in Urdu language and find that subjectivity in interpreting empathy among annotators appears to be independent of cultural background. Our systematic exploration of LMs’ understanding of empathy reveals substantial opportunities for further investigation in both task formulation and modeling.
@@ 1 Introduction
With Large Language Models (LLMs) demonstrating impressive capabilities in generating naturally sounding answers over a broad range of human inquiries, more individuals turn to seek solutions and emotional comfort by interacting with LLM-supported chatbots OpenAI (2023); Chang et al. (2024). They express thoughts, feelings and share their experiences, expecting deep understanding and sympathetic responses from a chatbot that can resonate with them, as shown in Figure 1 (Berridge et al., 2023). This requires LLMs to first fully understand the event, the emotions, and the empathy in the narratives, and then to respond appropriately Lin et al. (2023).
@@
Shen et al. (2023) proposed the task of measuring empathic similarity which assesses the similarity between two narratives describing personal experiences across four aspects: event, emotion, moral, and empathy, using a numerical score ranging from 1 to 4 (see more in Section 3.1). However, both fine-tuned LMs (BERT, RoBERTa) and few-shot prompted LLMs achieved low correlation with human annotations. Error analysis on Shen et al. (2023) shows that models can recognize dissimilar story pairs with scores in range of 1-2, but struggle to distinguish fine-grained differences between similar story pairs in range of 2.5-4 (see Section 3). This suggests that the nuanced patterns and relationships in similar pairs are not captured by current methods Wang et al. (2022a).
We hypothesize that the LMs used in these methods were primarily trained for semantic understanding tasks, rather than for emotion and empathy. They may capture the similarity between events, but they struggle to understand the complex social and emotional signals in human narratives Reimers and Gurevych (2019), this is further discussed in following sections. Moreover, naïve LLM prompts may not fully empower LLMs’ reasoning ability to make correct judgements. To this end, we explore multiple strategies to improve the accuracy of empathic similarity, including contrastive learning of LMs, LLM reasoning and fine-tuning with and without Chain-of-Thought (CoT) (Wei et al., 2022). We achieve around 5–10% improvement in the Pearson and Spearman correlations. However, accuracy appears to be capped and cannot be improved beyond a certain value.
This motivated us to speculate that gold labels might be problematic due to the subjective nature of judging emotion and empathy similarity. To investigate this, we randomly sampled ten story pairs from EmpathicStories and asked eight annotators with different backgrounds to annotate the similarity in terms of event, emotion, moral and empathy (defined in Section 3.1). Overall, low human agreement for all aspects was observed, especially moral, followed by empathy, emotion; the highest agreement was for the event. Interestingly, annotators from the same country/culture, particularly who are friends, had much higher agreement than others. We further collected a new Urdu dataset to explore the cultural impact on empathy similarity labeling, which revealed similar findings.
@@
  • We collect a new Urdu dataset to investigate the impact of language and culture on empathic labeling. Our analysis shows that subjectivity in the labeling is independent of the cultural background of annotators. [^1] Urdu dataset with individual similarity scores and modeling code are available at https://github.com/yuxiaw/Empathic-Similarity
2 Related Work
Emotion and Empathy in LMs:
Many prior work focus on tasks of recognizing (mixed) emotions, scoring intensity, inferring and explaining person’s emotional reactions (Liu et al., 2024b; Chen et al., 2024). In general domains, recent study illustrates that current LLMs are far better than human in generation, but fall short of understanding (West et al., 2023). However, for empathy-demanding scenarios, understanding and resonating with support-seekers are even more important than outputting opinions (i.e., generation) Buecker et al. (2021). In this work, we aim to examine language models capability in understanding and identifying nuanced difference of emotion and empathy, based on the task of empathic similarity (Shen et al., 2023). We aim to improve the estimation accuracy by forcing models to discern and learn the underlying reasons for similarities between pairs of stories, via (1) enhancing LM-based sentence embeddings, and (2) empowering LLM reasoning capability.
Sentence Embedding Enhancement:
Many previous methods improve semantic embeddings by adjusting loss functions in training Khan et al. (2022). SimCSE leverages contrastive loss Gao et al. (2021), and ESimCSE applies momentum contrast strategy to increase the number of negative pairs involved in the loss calculation, showing notable improvements across multiple semantic textual similarity benchmarks (Wu et al., 2021). Hubert et al. (2024) improved the knowledge graph embeddings by integrating semantic awareness into traditional loss functions like hinge loss. Huang et al. (2024) use a multi-task hybrid loss, incorporating both skeletal and semantic embedding into loss functions for micro-gesture classification. Therefore, we investigate a variety of loss functions such as ContrastiveLoss, CosineSimilarityLoss, CoSENTLoss and AnglELoss, based on LMs including MPNet Song et al. (2020), RoBERTa (base, large) Liu et al. (2019), DeBERTa (small, large) He et al. (2020) to improve sentence embedding in representing empathic features.
LLM Reasoning and Fine-tuning:
Naive LLM prompts may fail to fully leverage LLMs’ reasoning capabilities, leading to poor accuracy on EmpathicStories dataset Shen et al. (2023). Recent advancements in zero- and few-shot prompting techniques can boost performance such as Chain-of-Thought prompting (CoT) Wei et al. (2022), Least-to-Most prompting Zhou et al. (2022), and search-based approaches like Tree-of-Thought (ToT) Yao et al. (2024). We prompt LLMs by CoT and also fine-tune LLMs with reasoning trajectories, encouraging LLMs to think over before making the final decision.
@@ 3.1 Task and Dataset
In this work, we focus on the task of measuring the empathic similarity between two narratives by a numerical score in the range of 1 to 4, with 1 representing totally dissimilar and 4 indicating extremely similar. Empathic similarity assesses how much the narrators of a pair of stories would empathize with one another, in which main event, emotion, and moral similarities of two stories are core features influencing the empathic similarity.
Event highlights the similarity of main events in two experiences, as people empathize more with experiences that are similar to their own. Emotional reaction refers to how people emotionally respond to the experience. Individuals may have different feelings to the same experience, e.g., alone vs. happy for staying at home on weekend. Moral in this context emphasizes a higher level meaning abstracted by readers from a story, i.e., takeaway.
We used the EmpathicStories dataset, which was created by Shen et al. (2023). It consists of 1,500 unique stories and 2,000 story pairs, split into 1,500 pairs for training, 100 for development, and 400 for testing. Appendix Figure E shows the sample from EmpathicStories that consists of pair of stories and similarity scores on four labels. Table 1 shows the distribution of train, development, and test sets across different similarity ranges. Each story has two versions, full and summary, in which the annotators assign labels based on the summary from four perspectives including event, emotion, moral, and overall empathy. Figure 2 shows the correlation between the four similarity scores. We can see that the moral similarity has the highest correlation with empathy, followed by event and emotion.
Evaluation Measurements:
Pearson correlation ($r$) and Spearman rank correlation ($\rho$) are used to measure the performance of systems for predicting emotional similarity. This measures the linear correlation between the model outputs, the human annotations and the degree of monotonicity under ranking, respectively. Mean square error (MSE) assesses the models’ ability to get close to the gold standard. In a discrete setting, following Shen et al. (2023), both the predicted scores and the gold labels are binned into two classes thresholding by 2.5 — label = 1 if score > 2.5; otherwise, it is 0. Then accuracy, precision, recall and macro-F1 are used.
@@ 3.2 Error Analysis
Baseline: We reproduced the results of Shen et al. (2023) by fine-tuning SBERT Reimers and Gurevych (2019) and BART-base Lewis et al. (2019) as shown in Table 9, given that prompting davinci-text-003 and GPT-3.5-turbo had lower accuracy. We obtained a Spearman correlation $\rho$ similar to Shen et al. (2023), but with a lower F1.
@@
We found that the model excels at identifying dissimilar story pairs, but struggles with similar pairs with nuanced differences (not immediately obvious). Specifically, the model exhibits lower error rates in the score range of 1–2, and higher error rates in the range of 2.5–4. Particularly for 4, both models have a 100% error rate on the test set. We could attribute this to exposure bias during training. However, the number of training examples scoring between 2.5 and 4 is actually much larger than the number of dissimilar examples. This suggests that it is difficult for models to discern the subtle differences between similar story pairs, underlining a critical area for model improvement.
Hypothesis:
We hypothesize that SBERT and BART are primarily pre-trained and fine-tuned to learn semantic features, in which emotional and empathetic features are under-represented. Despite the importance of understanding the semantics of the stories, this does not necessarily account for capturing the deep, empathetic connections between narratives, especially when these connections are subtler and more implicit than straightforward lexical or thematic commonalities. The failure to identify the connections between similar narratives in terms of empathy leads to poor performance on similar story pairs ranging from 2.5 to 4, tending to be recognized as dissimilar with a score between 1-2. To alleviate these issues and to enhance the model performance on the empathetic similarity task, we propose a variety of strategies in Section 4.
@@ 4 Enhancing Empathy Representation
We explored various contrastive losses, including the ContrastiveLoss (Hadsell et al., 2006), AnglELoss (Li and Li, 2024) and CoSENTLoss333www.sbert.net. We also re-implemented the approach using cosine similarity loss as a baseline. We used masked language models (MLMs) including MPNet (Song et al., 2020), RoBERTa (base, large) (Liu et al., 2019) as text encoders leveraging their bidirectional encoding capabilities and computational efficiency.
LLM Reasoning and Fine-tuning:
@@
For the CoT explanation, even though the human-annotated explanations were provided in EmpathicStories, they were collected based on the story summary by GPT-3.5-Turbo, which can be biased and inaccurate. Moreover, the distribution of human-written text is distinct from the distribution of LLM-generated reasons, which makes the reasons provided by the humans less useful for lowering the perplexity in LLM fine-tuning  (Liu et al., 2024a). To address this, we prompted an LLM to generate explanations guided by gold labels, leveraging LLM capability of causal reasoning. That is, to explain why two narratives have the similarity of [scores placeholder] from the aspects of event, emotion, moral, and empathy.
Gold Label-Guided Explanation:
Based on the pair of full stories with the ground truth similarity scores, we asked the LLM to analyze the story pair from dimensions such as theme, emotional content, characters, narrative structures, and overall empathy. The generated analysis served as the reasoning content in supervised fine-tuning (SFT). We explored both full parameter and PEFT Hu et al. (2021) fine-tuning. We used Llama3-70B-instruct to generate explanations by prompts in Figure 4 of Appendix B.1.
@@ 5.1 Discriminative LMs
We refer to discriminative models as smaller pre-trained LMs like RoBERTa, and sentence embedding models such as SBERT. In both zero-shot and fine-tuning settings, we calculated cosine similarity using the embeddings of two stories generated by either pre-trained or fine-tuned sentence encoders.
@@ 5.1.2 Fine-Tuning with Contrastive Loss
We examined the effectiveness of contrastive losses presented in Section 4 based on three LMs. We fine-tuned them on the four types of gold labels separately and evaluated them using the overall empathy score as ground-truth based on pairs of story summary. [^4] The full story always exceeds the maximum input length of these three LMs, resulting in poor performance. This enables us to assess the impact of each attribute on modeling empathic similarity.
@@ 5.2.1 Zero-Shot Evaluation
Table 3 shows two major findings: (i) Zero-shot LLMs have clear advantages over cosine similarity based on sentence embeddings. $r$=0.442 for GPT-4o vs. $r$=0.336 for OpenAI-text-embedding on story summary. The larger the model, the higher the correlations with the gold empathy scores. This suggests that LLMs benefit from large-scale pre-training that empowers the distinguishability of nuanced empathic differences in narratives. (ii) Over all models and over the four types of scores, using the summaries is superior to using the full stories, which aligns with the fact that the gold labels are annotated based on the summary.
@@ 5.2.2 Is SFT Helpful?
To further investigate the potential of LLMs in this task, we performed supervised fine-tuning (SFT) with LLaMA-3-8B using the story summaries and the corresponding empathic similarity scores from the training set. In our experiments, we explored two different prompting strategies as introduced in Section 4. For the Score-Only strategy, we implemented both full-parameter and LoRA (Hu et al., 2021) SFT. For the CoT strategy, we only performed LoRA SFT due to the memory pressure of the longer sequence length. For all strategies, we tuned the model for two epochs to avoid overfitting.
@@ 5.3 Understanding the Bottleneck
To understand what LLMs learned during fine-tuning and why they struggle with this task, we analyzed the predictions made by the fully fine-tuned model. We selected this model because it is specifically adapted to the task of empathic similarity estimation, minimizing interference from its pre-trained knowledge.
@@
Regardless of the similarity class, the predicted probability kept the same as the empirical distribution of training data over four classes, i.e., $P(Y)$ = (0.140, 0.399, 0.404, 0.057), which is calculated by counting the percentage of story pairs falling into each class. That is, after fine-tuning, the model learned nothing about how to map the input $x\rightarrow y$, but merely the distribution of $P(Y)$. During inference, the model could not estimate the corresponding score conditioned on the input story pair, but sampled a similarity class based on the distribution of $P(Y)$ whatever the input pair was, leading to randomness on the development and the test sets (see confusion matrix of training set in Figure 5).
Based on these observations and previous findings in semantic textual similarity that models would be confused when training with ambiguous and subjective labels (Wang et al., 2022b, 2023, 2024), we speculate that the annotated empathic similarity scores are substantially subjective, i.e., that labels of high human disagreements hinder the model to learn distinguishable patterns. This could also explain why LMs and LLMs could not improve the performance by a large margin despite our extensive efforts. The struggle of the models to overcome the bottleneck can be attributed to the complexity and the variability in the human interpretations of these abstract concepts. This underscores the need for a critical analysis of the dataset’s quality and the inherent subjectivity of its labels.